# Complex exponential integral means spectrums of univalent functions and the Brennan conjecture

### v3

@@ 6.1. Univalent rational functions and the main result of this section
$({\bf c_{1}})$ if $R$ belongs to the class $\mathcal{R}_{O}\cap\mathcal{L}_{II}$, then
@@
In particular, we see from Corollary 6.20 that $\beta_{f}(\tau)\leq|\tau|-1$ for all $f\in\mathcal{U}_{R}$ when $\tau\leq-2$, which implies that
@@
(1) there are two constants $c_{0}>0$, $r_{0}\in(0,1)$ such that $|g(z)|\geq c_{0}$ for all $|z|\in(r_{0},1)$, and
@@
and $\Omega_{j_{1}}\cap\Omega_{j_{2}}=\emptyset$ whenever $j_{1}\neq j_{2}$.
@@
$\leq$ ${c_{0}}^{-\varkappa}\int_{-\pi}^{\pi}\Big[\sum_{j=1}^{n}\frac{|A_{j}|}{|re^{i\theta}-e^{i\arg z_{j}}|}\Big]^{\varkappa}d\theta$
$\leq$ ${c_{0}}^{-\varkappa}n^{\varkappa-1}{\mathbb{A}}^{\varkappa}\int_{-\pi}^{\pi}\sum_{j=1}^{n}\frac{d\theta}{|re^{i\theta}-e^{i\arg z_{j}}|^{\varkappa}}$
@@
$({\dagger})$ If $R$ has at east one critical point in $\mathbb{T}$, we let, as in (6.4),
@@
(1.1) there are always two constants $c_{0}>0$, $r_{0}\in(0,1)$ such that $|P_{\bf n}(z)/{P_{\bf m}(z)}|\geq c_{0}$ for all $|z|\in(r_{0},1)$, and
(1.2) for each $z_{j}=e^{i\arg z_{j}}$, we can always take constants $r_{j}\in(0,1)$, $\vartheta_{j}>0$, $C_{j}>0$ such that $|P_{\bf n}(z)/{P_{\bf m}(z)}|\leq C_{j}$ for all
$z\in\Omega_{j}:=\{z=re^{i\theta}:r_{j}<r<1,\arg z_{j}-\vartheta_{j}<\theta<\arg z_{j}+\vartheta_{j}\},\,\,j=1,2,\cdots,s,$
and $\Omega_{j_{1}}\cap\Omega_{j_{2}}=\emptyset$, whenever $j_{1}\neq j_{2}$. Then by using Lemma 6.24, we obtain that, as $r\rightarrow 1^{-}$,
$\int_{-\pi}^{\pi}\frac{d\theta}{|R^{\prime}(re^{i\theta})|^{\varkappa}}\asymp\begin{cases}\frac{1}{(1-r)^{\varkappa-1}},\;\text{if}\;\varkappa>1,\\ \log\frac{1}{1-r},\;\;\;\text{if}\;\varkappa=1,\\ 1,\quad\quad\quad\;\,\text{if}\;\varkappa\in(0,1).\end{cases}$
This implies that
@@
$({\ddagger})$ If $R$ has no critical points in $\mathbb{T}$. Then, it is easy to see from the definition (6.7) of the generalized integral means spectrum that ${\widetilde{\beta}}_{R}(\tau)=0$ for any $\tau<0$.
(C) We next consider the case $\tau>0$. Our proof contains the following three cases.
Case 1. When $R$ belongs to the class $\mathcal{R}_{O}\cap\mathcal{L}_{I}$, we easily see that ${\widetilde{\beta}}_{R}(\tau)=0$ for any $\tau>0$.
Case 2. When $R$ belongs to the class $\mathcal{R}_{O}\cap\mathcal{L}_{II}$. First we note that there is a constant $r_{R}\in(0,1)$ such that $R^{\prime}(z)\neq 0$ for all $z\in\{z=re^{i\theta}:r_{R}<r<1\}$. We write
@@
$(\bf 2.1)$ $\Omega_{j_{1}}\cap\Omega_{j_{2}}=\emptyset$ whenever $j_{1}\neq j_{2}$, $1\leq j_{1},j_{2}\leq l$, and $M_{j,2}\leq|Q_{1}(z)|\leq M_{j,1}$ for all $z\in\Omega_{j}$, here $\Omega_{j}=\{re^{\theta}:r_{0}<r<1,\theta_{j}-\eta_{j}\leq\theta\leq\theta_{j}+\eta_{j}\},$ $j=1,2,\cdots,l$, and
$(\bf 2.2)$ $|R^{\prime}(z)|\leq M$ for all $z\in\Omega_{0}-\bigcup\limits_{1\leq j\leq l}\Omega_{j}$, here $\Omega_{0}=\{re^{\theta}:r_{0}<r<1\}.$
@@
$D^{j}:=\sup\limits_{z\in\Omega_{j}}\{\prod\limits_{1\leq a\leq l,a\neq j}|z-e^{i\theta_{a}}|\},\,\,D_{j}:=\inf\limits_{z\in\Omega_{j}}\{\prod\limits_{1\leq a\leq l,a\neq j}|z-e^{i\theta_{a}}|\}.$
@@
It is easy to see that $0<\mathbf{D}_{\ast}<\mathbf{D}^{\ast}<+\infty$. Then, when $r>r_{0}$, we have
$\int_{-\pi}^{\pi}|R^{\prime}(re^{i\theta})|^{\tau}d\theta$ $\geq$ $M_{2}[\mathbf{D}^{*}]^{-2\tau}\sum_{j=1}^{l}\int_{\theta_{j}-\eta_{j}}^{\theta_{j}+\eta_{j}}\frac{d\theta}{|re^{i\theta}-e^{i\theta_{j}}|^{2\tau}}$
$\geq$ $M_{2}[\mathbf{D}^{*}]^{-2\tau}\sum_{j=1}^{l}\int_{-\eta_{j}}^{\eta_{j}}\frac{d\theta}{|1-re^{i\theta}|^{2\tau}},$
@@
$\int_{-\pi}^{\pi}|R^{\prime}(re^{i\theta})|^{\tau}d\theta$ $\leq$ $\sum\limits_{j=1}^{l}\int_{\theta_{j}-\eta_{j}}^{\theta_{j}+\eta_{j}}|R^{\prime}(re^{i\theta})|^{\tau}d\theta+2\pi M$
$\leq$ $M_{1}[{\bf{D}}_{*}]^{-2\tau}\sum\limits_{j=1}^{l}\int_{\theta_{j}-\eta_{j}}^{\theta_{j}+\eta_{j}}\frac{d\theta}{|re^{i\theta}-e^{i\theta_{j}}|^{2\tau}}+2\pi M$
$=$ $M_{1}[{\bf{D}}_{*}]^{-2\tau}\sum\limits_{j=1}^{l}\int_{-\eta_{j}}^{\eta_{j}}\frac{d\theta}{|1-re^{i\theta}|^{2\tau}}+2\pi M,$
@@
and for $\tau\in(0,\frac{1}{2}]$,
$\int_{-\pi}^{\pi}|R^{\prime}(re^{i\theta})|^{\tau}d\theta\asymp 1,\text{as}\,\,r\rightarrow 1^{-}.$
@@
Case 3. When $R$ belongs to the class $\mathcal{R}_{O}\cap\mathcal{L}_{III}$. First we note that there is a constant $r_{R}\in(0,1)$ such that $R^{\prime}(z)\neq 0$ for all $z\in\{z=re^{i\theta}:r_{R}<r<1\}$. We write
@@
$(\bf 3.1)$ $\Omega_{j_{1}}\cap\Omega_{j_{2}}=\emptyset$ whenever $j_{1}\neq j_{2}$ and $M_{j,2}\leq|Q_{2}(z)|\leq M_{j,1}$ for all $z\in\Omega_{j}$, here $\Omega_{j}=\{re^{\theta}:r_{0}<r<1,\theta_{j}-\eta_{j}\leq\theta\leq\theta_{j}+\eta_{j}\},$ $1\leq j_{1},j_{2}\leq l,$ and
$(\bf 3.2)$ $\widetilde{\Omega}_{k_{1}}\cap\widetilde{\Omega}_{k_{2}}=\emptyset$ whenever $k_{1}\neq k_{2}$, $1\leq k_{1},k_{2}\leq t$, and ${\Omega}_{j}\cap\widetilde{\Omega}_{k}=\emptyset$ for all $1\leq j\leq l,1\leq k\leq t$, and $\widetilde{M}_{k,2}\leq|Q_{2}(z)|\leq\widetilde{M}_{k,1}$ for all $z\in\widetilde{\Omega}_{k}$, here $\widetilde{\Omega}_{k}=\{re^{\theta}:{r}_{0}<r<1,\widetilde{\theta}_{k}-\widetilde{\eta}_{k}\leq\theta\leq\widetilde{\theta}_{k}+\widetilde{\eta}_{k}\},$ and
$(\bf 3.3)$ $|R^{\prime}(z)|\leq M$ for all $z\in\Omega_{0}-\bigcup\limits_{1\leq j\leq l,1\leq k\leq t}(\Omega_{j}\bigcup\widetilde{\Omega}_{k})$, here $\Omega_{0}=\{re^{\theta}:r_{0}<r<1\}.$
For $1\leq j\leq l$, we let $\widetilde{\pi}(z):=\prod\limits_{1\leq k\leq t}|z-e^{i\widetilde{\theta}_{k}}|^{2\tau}\}$ and take
$D^{j}:=\sup\limits_{z\in\Omega_{j}}\{\widetilde{\pi}(z)\prod\limits_{1\leq a\leq l,a\neq j}|z-e^{i\theta_{a}}|^{3\tau}\},\,\,D_{j}:=\inf\limits_{z\in\Omega_{j}}\{\widetilde{\pi}(z)\prod\limits_{1\leq a\leq l,a\neq j}|z-e^{i\theta_{a}}|^{3\tau}\}.$
@@
For $1\leq k\leq t$, we let $\pi(z):=\prod\limits_{1\leq j\leq l}|z-e^{i{\theta}_{j}}|^{3\tau}$ and take
$\widetilde{D}^{k}:=\sup\limits_{z\in\widetilde{\Omega}_{k}}\{\pi(z)\prod\limits_{1\leq b\leq t,b\neq k}|z-e^{i\widetilde{\theta}_{b}}|^{2\tau}\},\,\,\widetilde{D}_{k}:=\inf\limits_{z\in\widetilde{\Omega}_{k}}\{\pi(z)\prod\limits_{1\leq b\leq t,b\neq k}|z-e^{i\widetilde{\theta}_{b}}|^{2\tau}\}.$
@@
$\mathbf{D}^{\ast}=\max\limits_{1\leq j\leq l,1\leq k\leq t}\{D^{j},\widetilde{D}^{k}\},\,\,\mathbf{D}_{\ast}=\min\limits_{1\leq j\leq l,1\leq k\leq t}\{D_{j},\widetilde{D}_{k}\}.$
We easily see that $0<\mathbf{D}_{\ast}<\mathbf{D}^{\ast}<+\infty$. Then, when $r>r_{0}$, we have
@@
$\geq M_{2}[\mathbf{D}^{\ast}]^{-1}\Big[\sum_{j=1}^{l}\int_{\theta_{j}-\eta_{j}}^{\theta_{j}+\eta_{j}}\frac{d\theta}{|re^{i\theta}-e^{i\theta_{j}}|^{3\tau}}+\sum_{k=1}^{t}\int_{\widetilde{\theta}_{k}-\widetilde{\eta}_{k}}^{\widetilde{\theta}_{k}+\widetilde{\eta}_{k}}\frac{d\theta}{|re^{i\theta}-e^{i\widetilde{\theta}_{k}}|^{2\tau}}\Big]$
$=M_{2}[\mathbf{D}^{\ast}]^{-1}\Big[\sum_{j=1}^{l}\int_{-\eta_{j}}^{\eta_{j}}\frac{d\theta}{|1-re^{i\theta}|^{3\tau}}+\sum_{k=1}^{t}\int_{-\widetilde{\eta}_{k}}^{\widetilde{\eta}_{k}}\frac{d\theta}{|1-re^{i\theta}|^{2\tau}}\Big],$
@@
$\int_{-\pi}^{\pi}|R^{\prime}(re^{i\theta})|^{\tau}d\theta\leq\sum\limits_{j=1}^{l}\int_{\theta_{j}-\eta_{j}}^{\theta_{j}+\eta_{j}}|R^{\prime}(re^{i\theta})|^{\tau}d\theta+\sum\limits_{k=1}^{t}\int_{\widetilde{\theta}_{k}-\widetilde{\eta}_{k}}^{\widetilde{\theta}_{k}+\widetilde{\eta}_{k}}|R^{\prime}(re^{i\theta})|^{\tau}d\theta+2\pi M$
$\leq M_{1}[\mathbf{D}_{*}]^{-1}\Big[\sum\limits_{j=1}^{l}\int_{\theta_{j}-\eta_{j}}^{\theta_{j}+\eta_{j}}\frac{d\theta}{|re^{i\theta}-e^{i\theta_{j}}|^{3\tau}}+\sum\limits_{k=1}^{t}\int_{\widetilde{\theta}_{k}-\widetilde{\eta}_{k}}^{\widetilde{\theta}_{k}+\widetilde{\eta}_{k}}\frac{d\theta}{|re^{i\theta}-e^{i\widetilde{\theta}_{k}}|^{2\tau}}\Big]+2\pi M$
$=M_{1}[\mathbf{D}_{*}]^{-1}\Big[\sum_{j=1}^{l}\int_{-\eta_{j}}^{\eta_{j}}\frac{d\theta}{|1-re^{i\theta}|^{3\tau}}+\sum_{k=1}^{t}\int_{-\widetilde{\eta}_{k}}^{\widetilde{\eta}_{k}}\frac{d\theta}{|1-re^{i\theta}|^{2\tau}}\Big]+2\pi M,$
@@
and for $\tau\in(0,\frac{1}{3}]$,
$\int_{-\pi}^{\pi}|R^{\prime}(re^{i\theta})|^{\tau}d\theta\asymp 1,\,\,\text{as}\,\,r\rightarrow 1^{-}.$
@@
For the remaining case when $t=0$, by using again (c) in Remark 6.16 and by the similar way as in Case 2, we still can prove that (6.22) holds. Finally, combining all above arguments, we conclude that Theorem 6.18 is true. This finishes the proof of Theorem 6.18. ∎

### v5

@@ 6.1. Univalent rational functions and the main result of this section
$({\bf c_{2}})$ if $R$ belongs to the class $\mathcal{R}_{O}\cap\mathcal{L}_{II}$, then
@@
In particular, we see from Corollary 6.20 that $\beta_{f}(\tau)={\widetilde{\beta}}_{f}(\tau)\leq|\tau|-1$ for all $f\in\mathcal{U}_{R}$ when $\tau\leq-2$, which implies that
@@
(1) there are two constants $c_{0}>0$, $r_{0}\in(0,1)$ such that $|g^{-1}(z)|\leq c_{0}$ for all $|z|\in(r_{0},1)$, and
@@
and $\Omega_{j_{1}}\cap\Omega_{j_{2}}=\emptyset$ whenever $j_{1}\neq j_{2}$, $1\leq j_{1},j_{2}\leq n$.
@@
$\leq$ ${c_{0}}^{\varkappa}\int_{-\pi}^{\pi}\Big[\sum_{j=1}^{n}\frac{|A_{j}|}{|re^{i\theta}-e^{i\arg z_{j}}|}\Big]^{\varkappa}d\theta$
$\leq$ ${c_{0}}^{\varkappa}n^{\varkappa-1}{\mathbb{A}}^{\varkappa}\int_{-\pi}^{\pi}\sum_{j=1}^{n}\frac{d\theta}{|re^{i\theta}-e^{i\arg z_{j}}|^{\varkappa}}$
@@
$({\bf b_{1}})$ If $R$ has at east one critical point in $\mathbb{T}$, we let, as in (6.4),
$R^{\prime}(z)=\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{P_{\bf m}(z)}:=\prod_{j=1}^{s}(z-z_{j})g(z).$
@@
(1.1) there are always two constants $c_{0}>0$, $r_{0}\in(0,1)$ such that $|g^{-1}(z)|\leq c_{0}$ for all $|z|\in(r_{0},1)$, and
(1.2) for each $z_{j}=e^{i\arg z_{j}}$, $j=1,2,\cdots,s$, we can always find constants $r_{j}\in(0,1)$, $\vartheta_{j}>0$, $C_{j}>0$ such that $|g(z)|\leq C_{j}$ for all
$z\in\Omega_{j}=\{z=re^{i\theta}:r_{j}\leq r\leq 1,\arg z_{j}-\vartheta_{j}\leq\theta\leq\arg z_{j}+\vartheta_{j}\},$
and $\Omega_{j_{1}}\cap\Omega_{j_{2}}=\emptyset$, whenever $j_{1}\neq j_{2}$, $1\leq j_{1},j_{2}\leq s$, and $g(z)=P_{\bf n}(z)/{P_{\bf m}(z)}$ has no poles in $\bigcup\limits_{1\leq j\leq s}\Omega_{j}$. Then by using Lemma 6.24, we obtain that, as $r\rightarrow 1^{-}$,
(6.22) $\int_{-\pi}^{\pi}\frac{d\theta}{|R^{\prime}(re^{i\theta})|^{-\tau}}\asymp\begin{cases}\frac{1}{(1-r)^{-\tau-1}},\;\text{if}\;\tau<-1,\\ \log\frac{1}{1-r},\;\;\;\;\;\text{if}\;\tau=-1,\\ 1,\quad\quad\quad\;\,\;\;\text{if}\;\tau\in(-1,0).\end{cases}$
Note that for any small $\epsilon>0$,
$\log\frac{1}{1-r}/\frac{1}{(1-r)^{\epsilon}}=(1-r)^{\epsilon}\log\frac{1}{1-r}\rightarrow 0,\,\,{\text{as}}\,\,r\rightarrow 1^{-}.$
This implies that ${\widetilde{\beta}}_{R}(-1)=0$. Then it follows from (6.22) that
@@
$({\bf b_{2}})$ If $R$ has no critical points in $\mathbb{T}$. Then, it is easy to see from the definition (6.7) of the generalized integral means spectrum that ${\widetilde{\beta}}_{R}(\tau)=0$ for any $\tau<0$.
(C) We next consider the case $\tau>0$.
$({\bf c_{1}})$ When $R$ belongs to the class $\mathcal{R}_{O}\cap\mathcal{L}_{I}$, we easily see that ${\widetilde{\beta}}_{R}(\tau)=0$ for any $\tau>0$.
$({\bf c_{2}})$ When $R$ belongs to the class $\mathcal{R}_{O}\cap\mathcal{L}_{II}$. First we note that there is a constant $r_{R}\in(0,1)$ such that $R^{\prime}(z)\neq 0$ for all $z\in\{z=re^{i\theta}:r_{R}<r<1\}$. We write
@@
$(\bf 2.1)$ $\Omega_{j_{1}}\cap\Omega_{j_{2}}=\emptyset$ whenever $j_{1}\neq j_{2}$, $1\leq j_{1},j_{2}\leq l$, and $M_{j,2}\leq|Q_{1}(z)|\leq M_{j,1}$ for all $z\in\Omega_{j}$, here $\Omega_{j}=\{re^{\theta}:r_{0}\leq r\leq 1,\theta_{j}-\eta_{j}\leq\theta\leq\theta_{j}+\eta_{j}\},$ $j=1,2,\cdots,l$, and $R$ has no critical points in $\bigcup\limits_{1\leq j\leq l}\Omega_{j}$, and
$(\bf 2.2)$ $|R^{\prime}(z)|\leq M$ for all $z\in\Omega_{0}-\bigcup\limits_{1\leq j\leq l}\Omega_{j}$, here $\Omega_{0}=\{re^{\theta}:r_{0}\leq r\leq 1\}.$
@@
$D^{j}:=\max\limits_{z\in\Omega_{j}}\{\prod\limits_{1\leq a\leq l,a\neq j}|z-e^{i\theta_{a}}|\},\,\,D_{j}:=\min\limits_{z\in\Omega_{j}}\{\prod\limits_{1\leq a\leq l,a\neq j}|z-e^{i\theta_{a}}|\}.$
@@
It is easy to see that $0<\mathbf{D}_{\ast}<\mathbf{D}^{\ast}<+\infty$ since $\Omega_{j},j=1,2,\cdots,l$ are all compact. Then, when $r>r_{0}$, we have
$\int_{-\pi}^{\pi}|R^{\prime}(re^{i\theta})|^{\tau}d\theta$ $\geq$ $\sum_{j=1}^{l}\int_{\theta_{j}-\eta_{j}}^{\theta_{j}+\eta_{j}}\frac{|Q_{1}(re^{i\theta})|^{\tau}}{\prod\limits_{j=1}^{l}|re^{i\theta}-e^{i\theta_{j}}|^{2\tau}}d\theta$
$\geq$ $M_{2}^{\tau}[\mathbf{D}^{*}]^{-2\tau}\sum_{j=1}^{l}\int_{\theta_{j}-\eta_{j}}^{\theta_{j}+\eta_{j}}\frac{d\theta}{|re^{i\theta}-e^{i\theta_{j}}|^{2\tau}}$
$=$ $M_{2}^{\tau}[\mathbf{D}^{*}]^{-2\tau}\sum_{j=1}^{l}\int_{-\eta_{j}}^{\eta_{j}}\frac{d\theta}{|1-re^{i\theta}|^{2\tau}},$
@@
$\int_{-\pi}^{\pi}|R^{\prime}(re^{i\theta})|^{\tau}d\theta$ $\leq$ $\sum\limits_{j=1}^{l}\int_{\theta_{j}-\eta_{j}}^{\theta_{j}+\eta_{j}}|R^{\prime}(re^{i\theta})|^{\tau}d\theta+2\pi M^{\tau}$
$\leq$ $M_{1}^{\tau}[{\bf{D}}_{*}]^{-2\tau}\sum\limits_{j=1}^{l}\int_{\theta_{j}-\eta_{j}}^{\theta_{j}+\eta_{j}}\frac{d\theta}{|re^{i\theta}-e^{i\theta_{j}}|^{2\tau}}+2\pi M^{\tau}$
$=$ $M_{1}^{\tau}[{\bf{D}}_{*}]^{-2\tau}\sum\limits_{j=1}^{l}\int_{-\eta_{j}}^{\eta_{j}}\frac{d\theta}{|1-re^{i\theta}|^{2\tau}}+2\pi M^{\tau},$
@@
and for $\tau\in(0,\frac{1}{2}]$, $\text{as}\,\,r\rightarrow 1^{-},$
$\int_{-\pi}^{\pi}|R^{\prime}(re^{i\theta})|^{\tau}d\theta\asymp\begin{cases}\log\frac{1}{1-r},\;\;\;\text{if}\;\tau=\frac{1}{2},\\ 1,\quad\quad\quad\;\,\text{if}\;\tau\in(0,\frac{1}{2}).\end{cases}$
@@
$({\bf c_{3}})$ When $R$ belongs to the class $\mathcal{R}_{O}\cap\mathcal{L}_{III}$. The main idea of the proof of this case is similar to the case $({\bf c_{2}})$. First we note that there is a constant $r_{R}\in(0,1)$ such that $R^{\prime}(z)\neq 0$ for all $z\in\{z=re^{i\theta}:r_{R}<r<1\}$. We write
@@
$(\bf 3.1)$ $\Omega_{j_{1}}\cap\Omega_{j_{2}}=\emptyset$ whenever $j_{1}\neq j_{2}$ and $M_{j,2}\leq|Q_{2}(z)|\leq M_{j,1}$ for all $z\in\Omega_{j}$, here $\Omega_{j}=\{re^{\theta}:r_{0}\leq r\leq 1,\theta_{j}-\eta_{j}\leq\theta\leq\theta_{j}+\eta_{j}\},$ $1\leq j_{1},j_{2}\leq l,$ and
$(\bf 3.2)$ $\widetilde{\Omega}_{k_{1}}\cap\widetilde{\Omega}_{k_{2}}=\emptyset$ whenever $k_{1}\neq k_{2}$, $1\leq k_{1},k_{2}\leq t$, and ${\Omega}_{j}\cap\widetilde{\Omega}_{k}=\emptyset$ for all $1\leq j\leq l,1\leq k\leq t$, and $\widetilde{M}_{k,2}\leq|Q_{2}(z)|\leq\widetilde{M}_{k,1}$ for all $z\in\widetilde{\Omega}_{k}$, here $\widetilde{\Omega}_{k}=\{re^{\theta}:{r}_{0}\leq r\leq 1,\widetilde{\theta}_{k}-\widetilde{\eta}_{k}\leq\theta\leq\widetilde{\theta}_{k}+\widetilde{\eta}_{k}\},$ and $R$ has no critical points in $\bigcup\limits_{1\leq j\leq l,1\leq k\leq t}(\Omega_{j}\bigcup\widetilde{\Omega}_{k})$, and
$(\bf 3.3)$ $|R^{\prime}(z)|\leq M$ for all $z\in\Omega_{0}-\bigcup\limits_{1\leq j\leq l,1\leq k\leq t}(\Omega_{j}\bigcup\widetilde{\Omega}_{k})$, here $\Omega_{0}=\{re^{\theta}:r_{0}\leq r\leq 1\}.$
For $1\leq j\leq l$, we let $\widetilde{\pi}(z):=\prod\limits_{1\leq k\leq t}|z-e^{i\widetilde{\theta}_{k}}|^{2}\}$ and take
$D^{j}:=\max\limits_{z\in\Omega_{j}}\{\widetilde{\pi}(z)\prod\limits_{1\leq a\leq l,a\neq j}|z-e^{i\theta_{a}}|^{3}\},\,\,D_{j}:=\min\limits_{z\in\Omega_{j}}\{\widetilde{\pi}(z)\prod\limits_{1\leq a\leq l,a\neq j}|z-e^{i\theta_{a}}|^{3}\}.$
@@
For $1\leq k\leq t$, we let $\pi(z):=\prod\limits_{1\leq j\leq l}|z-e^{i{\theta}_{j}}|^{3}$ and take
$\widetilde{D}^{k}:=\max\limits_{z\in\widetilde{\Omega}_{k}}\{\pi(z)\prod\limits_{1\leq b\leq t,b\neq k}|z-e^{i\widetilde{\theta}_{b}}|^{2}\},\,\,\widetilde{D}_{k}:=\min\limits_{z\in\widetilde{\Omega}_{k}}\{\pi(z)\prod\limits_{1\leq b\leq t,b\neq k}|z-e^{i\widetilde{\theta}_{b}}|^{2}\}.$
@@
$\mathbf{D}^{\star}=\max\limits_{1\leq j\leq l,1\leq k\leq t}\{D^{j},\widetilde{D}^{k}\},\,\,\mathbf{D}_{\star}=\min\limits_{1\leq j\leq l,1\leq k\leq t}\{D_{j},\widetilde{D}_{k}\}.$
We easily see that $0<\mathbf{D}_{\star}<\mathbf{D}^{\star}<+\infty$ from the compactness of $\Omega_{j},j=1,2,\cdots,l$, and $\widetilde{\Omega}_{k},k=1,2,\cdots,t$. Then, when $r>r_{0}$, we have
@@
$\geq M_{2}^{\tau}[\mathbf{D}^{\star}]^{-\tau}\Big[\sum_{j=1}^{l}\int_{\theta_{j}-\eta_{j}}^{\theta_{j}+\eta_{j}}\frac{d\theta}{|re^{i\theta}-e^{i\theta_{j}}|^{3\tau}}+\sum_{k=1}^{t}\int_{\widetilde{\theta}_{k}-\widetilde{\eta}_{k}}^{\widetilde{\theta}_{k}+\widetilde{\eta}_{k}}\frac{d\theta}{|re^{i\theta}-e^{i\widetilde{\theta}_{k}}|^{2\tau}}\Big]$
$=M_{2}^{\tau}[\mathbf{D}^{\star}]^{-\tau}\Big[\sum_{j=1}^{l}\int_{-\eta_{j}}^{\eta_{j}}\frac{d\theta}{|1-re^{i\theta}|^{3\tau}}+\sum_{k=1}^{t}\int_{-\widetilde{\eta}_{k}}^{\widetilde{\eta}_{k}}\frac{d\theta}{|1-re^{i\theta}|^{2\tau}}\Big],$
@@
$\int_{-\pi}^{\pi}|R^{\prime}(re^{i\theta})|^{\tau}d\theta\leq\sum\limits_{j=1}^{l}\int_{\theta_{j}-\eta_{j}}^{\theta_{j}+\eta_{j}}|R^{\prime}(re^{i\theta})|^{\tau}d\theta+\sum\limits_{k=1}^{t}\int_{\widetilde{\theta}_{k}-\widetilde{\eta}_{k}}^{\widetilde{\theta}_{k}+\widetilde{\eta}_{k}}|R^{\prime}(re^{i\theta})|^{\tau}d\theta+2\pi M^{\tau}$
$\leq M_{1}^{\tau}[\mathbf{D}_{\star}]^{-\tau}\Big[\sum\limits_{j=1}^{l}\int_{\theta_{j}-\eta_{j}}^{\theta_{j}+\eta_{j}}\frac{d\theta}{|re^{i\theta}-e^{i\theta_{j}}|^{3\tau}}+\sum\limits_{k=1}^{t}\int_{\widetilde{\theta}_{k}-\widetilde{\eta}_{k}}^{\widetilde{\theta}_{k}+\widetilde{\eta}_{k}}\frac{d\theta}{|re^{i\theta}-e^{i\widetilde{\theta}_{k}}|^{2\tau}}\Big]+2\pi M^{\tau}$
$=M_{1}^{\tau}[\mathbf{D}_{\star}]^{-\tau}\Big[\sum_{j=1}^{l}\int_{-\eta_{j}}^{\eta_{j}}\frac{d\theta}{|1-re^{i\theta}|^{3\tau}}+\sum_{k=1}^{t}\int_{-\widetilde{\eta}_{k}}^{\widetilde{\eta}_{k}}\frac{d\theta}{|1-re^{i\theta}|^{2\tau}}\Big]+2\pi M^{\tau},$
@@
and for $\tau\in(0,\frac{1}{3}]$, $\text{as}\,\,r\rightarrow 1^{-},$
$\int_{-\pi}^{\pi}|R^{\prime}(re^{i\theta})|^{\tau}d\theta\asymp\begin{cases}\log\frac{1}{1-r},\;\;\;\text{if}\;\tau=\frac{1}{3},\\ 1,\quad\quad\quad\;\,\text{if}\;\tau\in(0,\frac{1}{3}).\end{cases}$
@@
For the remaining case when $t=0$, by the similar way as in the case $({\bf c_{2}})$ and using again (c) in Remark 6.16, we still can prove that (6.23) holds. Finally, combining all above arguments, we conclude that Theorem 6.18 is true. This finishes the proof of Theorem 6.18. ∎